OpenStudy (anonymous):

\[ y\ddot{y}-\dot{y}^2=1 \] Which method would I use here? The boundary conditions are \[y(a)=y(-a)=1 \]

OpenStudy (anonymous):

i think reduction of order with\[p=\dot y\]will lead us to something maybe?

OpenStudy (anonymous):

just note that\[\ddot y=p\dot p \]

OpenStudy (anonymous):

\[ \dot{p}=\frac{ dp }{ dx}= \frac{d^2y}{dx^2}=\ddot{y} \]Surely?

OpenStudy (anonymous):

[yy']'=yy''+y'^2 so just difference in sign. Maybe it is a derivative of some quotient.

OpenStudy (anonymous):

make p a function of y\[\ddot y=\frac{dp}{dx}=\frac{dp}{dy}\frac{dy}{dx}=p \dot p\] and what @myko mentioned can be a good start too

OpenStudy (anonymous):

actually that is\[\ddot y=\frac{dp}{dx}=\frac{dp}{dy}\frac{dy}{dx}=p \dot p_y\]

OpenStudy (anonymous):

Sorry, I got confused by the notation

OpenStudy (anonymous):

\[\Large yy'' - \left( y' \right)^2 = -\left(y'\right)^2\left( \frac{y}{y'} \right)'\]

OpenStudy (anonymous):

Very nice, but how does that help?